OpenStudy (koalagurl1):

I need to do a presentation on algorithms and I can't think of a good algorithm. Will fan and medal. Thanks! :)

OpenStudy (amistre64):

eulers algorithm for finding the GCD of 2 numbers is fairly interesting and used as a fundamental concept in number theory i believe

OpenStudy (koalagurl1):

Do you know any simple algorithms? @amistre64

OpenStudy (amistre64):

GCD is a simple algorithm

OpenStudy (koalagurl1):

ok

OpenStudy (amistre64):

given 2 number a,b such that a>b a= bn + r1 for some integer n b = r1n + r2 r1 = r2n + r3 .... when rk = 0, then the GCD is r{k-1} ------------------------- try 12 and 28 28 = 12(2) + 4 12 = 4(3) + 0 4 is the GCD of 12 and 18

OpenStudy (amistre64):

the GCD of 12 and 28 that is

OpenStudy (koalagurl1):

thanks! :)